# GPU-Accelerated and Scalable Optimization (ScaleOpt)

### v1

@@ 1 Introduction
Recent research underscores the need for integrated solutions that jointly optimize algorithmic and hardware utilization. Works such as GPU-accelerated sparse matrix multiplication mishra2021accelerating ; hoefler2021sparsity demonstrate that carefully tuned low-level kernels can yield substantial performance gains in matrix-heavy workloads. Similarly, distributed parallel approaches to computation illustrate how algorithmic restructuring can unlock hardware scalability poshtkohi2023implementing . Recent academic contributions span a wide range of areas, from physics-informed neural networks BAZMARA2023152 ; shaeri2025multimodalphysicsinformedneuralnetwork ; afzal2025physics ; firoozsalari2024machine to applications such as environmental modeling alkhaled2024webmrt ; karimi2023level ; shaeri2025sentimentsocialsignalsclimate , to multimodal fact-checking beigi2024can ; shaeri2023semi , dataset generation frameworks shaeri2025mnistgenmodularmniststyledataset , resource-efficient training methods azghan2025can ; azghan2024cudle , and efficient dual-path dropout layers shaeri2025midlmatrixinterpolateddropoutlayer . Across these diverse efforts, a common challenge emerges: current training pipelines operate under fixed memory and run-time quotas, which limit scalability and adaptability.

### v2

@@ 1 Introduction
Recent research underscores the need for integrated solutions that jointly optimize algorithmic and hardware utilization. Works such as GPU-accelerated sparse matrix multiplication mishra2021accelerating ; hoefler2021sparsity demonstrate that carefully tuned low-level kernels can yield substantial performance gains in matrix-heavy workloads. Similarly, distributed parallel approaches to computation illustrate how algorithmic restructuring can unlock hardware scalability poshtkohi2023implementing . Recent academic contributions span a wide range of areas, from physics-informed neural networks BAZMARA2023152 ; shaeri2025multimodal ; afzal2025physics ; firoozsalari2024machine to applications such as environmental modeling alkhaled2024webmrt ; karimi2023level ; shaeri2025sentiment , to multimodal fact-checking beigi2024can ; shaeri2023semi , dataset generation frameworks shaeri2025mnist ; jeong2025fediversesharing , resource-efficient training methods azghan2025can ; azghan2024cudle , and efficient dual-path dropout layers shaeri2025mid . Across these diverse efforts, a common challenge emerges: current training pipelines operate under fixed memory and run-time quotas, which limit scalability and adaptability.